# Huan Zhang

### v1

@@ 5 Numerical experiments
$\left\{\begin{array}[]{ll}&\mathop{\mbox{min}}\limits_{x\in\mathbb{R}^{n}}\|x\|_{1}+\frac{\mu}{2}\|x\|^{2}\\ &\mbox{s.t.}~~Ax=b,\end{array}\right$. (42)
@@
or the number of iterations exceeds 150. Here, the $z^{\prime}_{k}$ are the iterates generated by FISTA with the accuracy $\gamma:=\{10^{-4}$, $10^{-6}$, $10^{-8}\}$, respectively.
@@ 6 Conclusion
For future work, we plan to explore accelerated algorithms from the time discretization of the same differential system incorporating Tikhonov regularization to solve convex-concave bilinear saddle point problems. Furthermore, the framework presented here can be extended to develop accelerated algorithms for separable convex optimization problems.

### v2

@@ 5 Numerical experiments
$\left\{\begin{split}&\mathop{\mbox{min}}\limits_{x\in\mathbb{R}^{n}}\|x\|_{1}+\frac{\mu}{2}\|x\|^{2}\\ &\mbox{s.t.}~~Ax=b,\end{split}\right$. (36)
@@
or the number of iterations exceeds 150. Here, the $z^{\prime}_{k}$ are the iterates generated by FISTA with the accuracy $\gamma:=\{10^{-4}$, $10^{-6}$, $10^{-8}\}$, respectively. As noted in Remark 4.2, the subproblem generally does not admit an explicit closed-form solution. In this example, the subproblem is solved approximately by the FISTA with a sufficiently small stopping tolerance. Although the theoretical convergence of IAPDA is established under the assumption that the subproblem is solved exactly, such a highly accurate approximate solution can be regarded as a small perturbation of the exact one. As shown similarly in s2011 ; He2025 ; he2026 , this mild approximation error does not affect the convergence properties under proper error control.
@@ 6 Conclusion
Despite these advances, the convergence analysis of trajectories generated by IAPDA remains an open challenge. In future work, we aim to address this limitation by reformulating the system and constructing an appropriate energy function, with the goal of rigorously establishing trajectory convergence. Moreover, the framework presented here can be extended to develop accelerated algorithms for separable convex optimization problems. We also plan to explore accelerated algorithms from the time discretization of the same differential system incorporating Tikhonov regularization to solve convex-concave bilinear saddle point problems.
Acknowledgements
The authors would like to express their sincere thanks to the anonymous reviewers for the constructive suggestions and comments.